Find ALL the complex solutions to z^10-8z^5+64=0.
Answer #1

zo_82464=0 ze=t test+64 =0 = 8√ 64-256 2 = 875-192 2 = 8+ 813 i 2 t = 441B3. t = 4(1+iB), 4(1-BB) = 8(+is@).(-) = 8 (COS() +